the length is twice the width and area is 18
i have to find length and width

Found 2 solutions by richwmiller, unlockmath:
Answer by richwmiller(17219)

w*2w=18
2w^2=18
w^2=9
w=3 (we can't have a negative side)

Answer by unlockmath(1688)

Hello,
Let's represent the width with x and the length with 2x. To find the area we multiply length times width like this:
x(2x)=18 rewritten it will look like:
2x^2=18 Now divide both sides by 2
x^2=9 Now square root each side gives us:
x=3
So the width is 3 and the length is 6
Check it out. Multiply 3 times 6 and gives us 18
RJ
Check out the new book at:
www.math-unlock.com